F oh ND equation of line passing through (a cos x, a sin x ) and (a cos y,a sin y)

Grade:11

1 Answers

Novashree Samantaray
6 years ago
equation of line can be given by
[a(siny-sinx)/a(cosy-cosx)] (X-acosx) = Y-asinx
= [2sin({y-x}/2)cos({x+y}/2) / -2sin({y-x}/2)sin({x+y}/2)] (X-acosx) = Y-asinx
{cos(x+y)/2} (X-acosx) = -{sin(x+y)/2} (Y-asinx)
[cos(x+y)/2]X + [sin(x+y)/2]Y = a(sinx.sin(x+y)/2 + cosx.cos(x+y)/2)
                                              = a(cos(x-(x+y)/2))
                                              = a(cos(x-y)/2)
that is also the equation of chord joining points lying on a circle X^2 + Y^2 = a^2